# Analog In-memory Training on General Non-ideal Resistive Elements: The Impact of Response Functions

### v4

@@ I.5 Proof of Corollary 1: Exact convergence of Residual Learning
$E^{\texttt{RL}}_{K}\geq$ $\ \Omega\left(\gamma^{3}\|W_{k}-W^{*}\|^{2}_{R(W_{k})}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$ (123)
$\geq$ $\ \Omega\left(\min\{R(W_{k})\}\gamma^{3}\|W_{k}-W^{*}\|^{2}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$
@@
Further, if $R^{\texttt{ASGD}}_{\min}:=\min_{k\in[K]}\min\{Q_{+}(W_{k})Q_{-}(W_{k})\}>0$, it holds that
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KR^{\texttt{ASGD}}_{\min}}.$ (127)
@@
Define the saturation vector $R(W_{k})\in{\mathbb{R}}^{D}$ by
$R(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (132)
@@
where the last inequality choose $\alpha=\frac{1}{2LF_{\max}}$. Further, if the degree of saturation is bounded, (128)–(133) implies that
$\frac{\alpha R^{\texttt{AGD}}_{\min}}{2}\|\nabla f(W_{k})\|^{2}\leq\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}_{R(W_{k})}\leq f(W_{k})-f(W_{k+1}).$ (136)
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2(f(W_{0})-f(W_{K+1}))F_{\max}}{\alpha KR^{\texttt{AGD}}_{\min}}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KR^{\texttt{AGD}}_{\min}}$ (137)
@@
RPU Configuration. AIHWKIT offers fine-grained simulations of the hardware imperfections, such as the IO noise, analog-digital conversion, and so on. They are specified by the resistive processing unit (RPU) configurations. Without other specifications, we use the configuration list in Table˜4. The experimental setup uses a specific I/O configuration, as detailed in the relevant table. The system’s input and output signal bounds are explicitly defined. Regarding signal quality, the setup employs no input noise but introduces additive Gaussian noise to the output signal, the statistical properties of which are precisely specified. Finally, the resolution of the digital conversion processes is set by distinct bit values for both the input (DAC) and output (ADC).
In addition, noise, bound, and update management techniques are used in [71]. A learnable scaling factor is applied after each analog layer and updated using SGD. For each gradient update step, if more than $\text{BL}=32$ pulses are desired, only BL pulses are fired.

### v5

@@ I.5 Proof of Corollary 1: Exact convergence of Residual Learning
$E^{\texttt{RL}}_{K}\geq$ $\ \Omega\left(\gamma^{3}\|W_{k}-W^{*}\|^{2}_{M(W_{k})}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$ (129)
$\geq$ $\ \Omega\left(\min\{M(W_{k})\}\gamma^{3}\|W_{k}-W^{*}\|^{2}+\gamma^{2}\left\|P_{k}-P^{*}(W_{k})\right\|^{2}\right)$
@@
Further, if $M^{\texttt{ASGD}}_{\min}:=\min_{k\in[K]}\min\{Q_{+}(W_{k})\odot Q_{-}(W_{k})\}>0$, it holds that
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KM^{\texttt{ASGD}}_{\min}}.$ (133)
@@
Define the saturation vector $M(W_{k})\in{\mathbb{R}}^{D}$ by
$M(W_{k}):=$ $\ {F(W_{k})^{\odot 2}-G(W_{k})^{\odot 2}}={(F(W_{k})+G(W_{k}))\odot(F(W_{k})-G(W_{k}))}$ (138)
@@
where the last inequality choose $\alpha=\frac{1}{2LF_{\max}}$. Further, given the response functions are bounded, (134)–(139) implies that there is a lower bound $M^{\texttt{AGD}}_{\min}$ such that
$\frac{\alpha M^{\texttt{AGD}}_{\min}}{2}\|\nabla f(W_{k})\|^{2}\leq\frac{\alpha}{2}\|\nabla f(W_{k})\|^{2}_{M(W_{k})}\leq f(W_{k})-f(W_{k+1}).$ (142)
@@
$\frac{1}{K}\sum_{k=0}^{K-1}\|\nabla f(W_{k})\|^{2}\leq\frac{2(f(W_{0})-f(W_{K+1}))F_{\max}}{\alpha KM^{\texttt{AGD}}_{\min}}\leq\frac{2L(f(W_{0})-f^{*})F_{\max}^{2}}{KM^{\texttt{AGD}}_{\min}}$ (143)
@@
RPU Configuration. AIHWKIT offers fine-grained simulations of the hardware imperfections, such as the IO noise, analog-digital conversion, and so on. They are specified by the resistive processing unit (RPU) configurations. Without other specifications, we use the configuration list in Table˜4. The experimental setup uses a specific I/O configuration, as detailed in the relevant table. The system’s input and output signal bounds are explicitly defined. Regarding signal quality, the setup employs no input noise but introduces additive Gaussian noise to the output signal, the statistical properties of which are precisely specified. Finally, the resolution of the digital conversion process is determined by distinct bit values for both the input (DAC) and the output (ADC).
In addition, noise, bound, and update management techniques are used [71]. A learnable scaling factor is applied after each analog layer and updated using SGD. For each gradient update step, if more than $\text{BL}=32$ pulses are desired, only BL pulses are fired.